OpenStudy (demonchild99):

i think the answer is a...but im not sure If the volume for 3-D polyhedra A is 400 cm3 and the volume for 3-D polyhedral B is 800 cm3, what is the percent of change from A to B? 200% 20% 2% 0.2%

jagr2713 (jagr2713):

coRRECT

OpenStudy (demonchild99):

thx!

jagr2713 (jagr2713):

Anytime

OpenStudy (demonchild99):

What is the volume of grain that could completely fill this silo, rounded to the nearest whole number? Use 22 over 7 for pi. 13,750 ft3 14,012 ft3 262 ft3 4,583 ft3 the hight is 175 and the radius is 5

OpenStudy (demonchild99):

i think its a

OpenStudy (demonchild99):

@jagr2713

jagr2713 (jagr2713):

give me a sec

jagr2713 (jagr2713):

i dont know

OpenStudy (demonchild99):

lol...ok

jagr2713 (jagr2713):

ohohoh i figured it out u still here

OpenStudy (demonchild99):

yea...internet kicked out

OpenStudy (demonchild99):

@jagr2713

jagr2713 (jagr2713):

sorry um give me a sec really busy

OpenStudy (demonchild99):

ok lol

jagr2713 (jagr2713):

volume of a sphere V=4/3pir^3, for 1/2 sphere it would be 2/3 pi r^3 for this it would be (2/3) (22/7)(125)=261.905 cubic feet. Now for the cylinder portion. V=pi r^2 h=(22/7) (25)(175)=13750 cu ft Total=14,011.9 rounding off 14,012. You should of got that when you plugged the numbers in and solved.